A train takes 7 hours to travel from Station A to Station B. If its speed is increased by 10km/h. It will take 1 hour less to complete the journey. Find the distance between the two stations.
Answer by mananth(16946)

let speed of train be x
let distance be y
time taken = y/x=7
speed increased by 10 km/h
(x+10)
time = y/(x+10)
y/x - y/(x+10) =1
7-y/(x+10)=1
y/(x+10)= 6
y=6x+60
but y=7x
7x=6x+60
x=60 distance between two stations